When a metal plate is exposed to a monochromatic beam of light of wavelength 400 nm, a negative potential of 1.1 V is needed to stop the photocurrent. Find the threshold wavelength for the metal.

Grade:10

1 Answers

Navjyot Kalra
9 years ago
Sol. λ = 400 nm = 400 * 10^-9 m V base 0 = 1.1 V hc/λ = hc/λ base 0 + ev base 0 ⇒ 6.63 * 10^-34 *3 * 10^8/400 * 10^-9 = 6.63 * 10^-34 *3 * 10^8/λ base 0 ⇒ 4.97 = 19.89 * 10^-26/λ base 0 + 1.76 ⇒ 19.89 * 10^-26/λ base 0 = 4.97 – 17.6 = 3.21 ⇒ λ base 0 = 19.89 * 10^-26/3.21 = 6.196 * 10^-7 m = 620 nm.